# G-Adaptivity: optimised graph-based mesh relocation for finite element methods

### v2

@@ 5 Experimental results
We evaluate G-adaptivity on three classical meshing problems in two-dimensions: an elliptic PDE (Poisson’s equation) in a variety of convex domains, a nonlinear time-evolution PDE (Burger’s equation), and the time dependent Navier-Stokes equations in a multiply connected domain. In the following we present the performance improvements obtained in terms of the FEM $L^{2}$-error reduction (cf. (3)) and compute time, using our novel approach for adaptive meshing on each of these problems. Full code to build the datasets and reproduce our results will be made available.
@@ 5.1 Experimental details
Our experimental pipeline consists of three parts: (i) we build datasets containing information about the PDE, FEM solution on a regular grid and Hessian approximation. (ii) Then we train either our model or the baseline to predict a relocated mesh on which we then make FE solve for the solution; (iii) finally, we compare this FEM solution to a reference solution (calculated on a fine reference mesh) and determine the change in $L2$ error over the original undeformed mesh, i.e. in the above notation we look at relative error reduction of $E(U_{\mathcal{M}_{\theta}})$ over $E(U_{\mathcal{Z}_{0}})$. These steps are repeated for the three PDE datasets as described below.
@@
For each experiment described below we build a randomised dataset with held-out test data. This means in each case we specify a set of solution values through varying source terms or boundary conditions (adjusted to the PDE at hand). We then generate training and test sets of coarse, undeformed, meshes $\mathcal{T}$ of varying resolution with associated FEM solution values and, for each mesh, we also compute a reference solution (on a finer reference mesh) which serves as comparison for the error computation. For most test cases the solution values are generated by randomly sampling Gaussians in the domain $\Omega$. For the final example, the Navier–Stokes flow around a cylinder we used snapshots of a time-series simulation of vortex shedding. Full details on the specific configurations for each experiment are provided in Appendix C.
@@
Here we refer to our framework G-Adaptivity and our model Diffformer synonymously, which we train using the regularized PDE-loss (9) (which is the regularized $L^{2}$ FEM approximation error + equidistribution regularizer). Calculation of the $L^{2}$-error is obtained by calculating an FE solution on the moved mesh and comparing this to the projection of the reference solution of a fine regular reference mesh onto sufficiently higher order elements. We train our new model (G-Adapt) and UM2N-G for 300 epochs using an Adam optimiser and learning rate of 0.001. Our model has $4$ diffformer blocks as described in Appendix C. Each blocked is rolled out using explicit Euler integration for 32 steps with a step size of 0.1. For the baseline UM2N we trained using 1000 epochs in order to achieve good performance but we believe further tuning of the training may be required in order to achieve a similar performance to the one reported in (Zhang et al., 2024). UM2N remains an important baseline and we expect that with appropriate training the method would be able to achieve a similar error reduction (ER) as the Monge-Ampère (MA) solver, but we would like to highlight that even in the best reported results of the original paper, UM2N never achieved an larger error reduction than MA.
@@
We report three metrics to evaluate the performance of the mesh relocation methods at hand: (i) the relative $L^{2}$ error reduction (ER) of the FE solution on the relocated mesh versus the FE solution on the initial coarse mesh (larger error reduction means improved performance); (ii) the time taken to relocate the mesh (shorter times means faster relocation); (iii) the aspect ratio of the deformed mesh as a measure of mesh quality as described in Appendix F.4 (a single digit aspect ratio generally acceptable and smaller aspect ratio means more regular mesh).Each experiment is performed against 5 random seeds for the training to provide the error bars.
@@ 5.2 Benchmarking on Poisson’s equation
Our first benchmark is on the classical Poisson problem $-\nabla^{2}u=f({\mathbf{z}})$ with Dirichlet boundary conditions. Full details of the FEM formulation are given in Appendix C.1. We benchmark the results against two datasets on a square (cf. Figure 1) and polygonal domain respectively (cf. Figure 4). For both examples we sample source terms and boundary conditions corresponding to underlying Gaussian fields. On the square domain $\Omega=[0,1]^{2}$ we initialise the mesh-deformation with a regular grid and to showcase G-adaptivty’s ability to work on irregular domains with unstructured meshes we apply a similar methodology to the convex polygonal dataset (a sample is shown in Figure 4). The results for both datasets are presented in Table 1. The central observation in these results is that our methodology provides the very first ML approach to mesh relocation which is able to outperform MA in terms of error reduction, while retaining the fast mesh relocation times given by the state-of-the-art GNN - UM2N (Zhang et al., 2024).
@@ 5.3 Time-dependent Burgers’ equation
We train the models on the initial Gaussian for a timestep $\delta t=0.02$ with 2 steps and evaluate by following 10 trajectories of randomly sampled Gaussians in the Burgers equation for 20 timesteps, remeshing after every 2 steps (cf. Figure 5). The results in the top part of Table 2 show the average error reduction over achieved over every block of two timesteps. While the MA performs well on this task, we note that the UM2N and UM2N-G baselines appear to lead to a negative error reduction (i.e. an increase), which is likely due to the fact that the Burgers’ equation changes the solution shape and thus trajectories will lead to out-of-distribution cases for methods that are trained only on initial conditions. Due to the structural regularity of our new approach (cf. Section 4.2) our approach is able to deal with out-of-distribution data very well, and most importantly is able to outperform MA in terms of error reduction while retaining a fast mesh relocation time.
@@ Acknowledgements
JR, TD, CBS & CB gratefully acknowledge support from the EPSRC programme grant in ‘The Mathematics of Deep Learning’, under the project EP/V026259/1. GM gratefully acknowledges funding from the Mathematical Institute, University of Oxford. KS gratefully acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 850941).
@@ A.3 Gaussian basis function expansion for time-dynamic training
Appendix B An introduction to Firedrake
@@
Appendix C Further details on the test problems and finite element formulations
@@ D.1 Model and data hyperparameters
D.2 Additional results
In Figures 7 and 8 we present additional plots from the Poisson experiments on the square and polygonal domain detailed in the main paper exhibiting the types of meshes generated with our novel G-Adaptive methodology.
@@ Appendix E Further details on mesh relocation
E.1 The Equidistribution Principle
@@
where $\theta$ (to be determined) is a constant. The function $m$ is usually a function of $u$. An important example is given by the problem of linearly interpolating $u({\mathbf{z}})$ as it follows from Céa’s lemma that the resulting interpolation error is (an often tight) upper bound for the FE solution error. In this case $m$ will be a function of the curvature of $u$ (with the exact form dependent on the norm used to measure the error) (Huang & Russell, 2011).
@@
In one dimension the equidistribution condition (24) uniquely defines each cell length, and thus the cell shape, and hence the whole mesh. However in two dimensions it only gives the cell area but not the shape. To find the mesh uniquely additional conditions must be imposed. Noting the correspondence between the equidistribution condition and a measure on the physical space, the deformation map can be viewed as mapping a uniform measure in the computational space to a new measure in the physical space. It is natural to seek a map which minimises the cost of doing this, as this leads to meshes in the physical domain which are close to uniform and hence have minimal skewness and which avoid tangling. This gives an obvious link between mesh generation and optimal transport. In the continuous setting such a map can be calculated by solving (either directly or by using a surrogate solver) an associated Monge-Ampére equation, leading to the MA methods described in the main body of the text. (Note that with modifications this procedure can also be used to generate meshes on non-planar manifolds (McRae et al., 2018).)

### v3

@@ 5 Experimental results
We evaluate G-adaptivity on three classical meshing problems in two-dimensions: an elliptic PDE (Poisson’s equation) in a variety of convex domains, a nonlinear time-evolution PDE (Burger’s equation), and the time dependent Navier-Stokes equations in a multiply connected domain. In the following we present the performance improvements obtained in terms of the FEM $L^{2}$-error reduction (cf. (3)) and compute time, using our novel approach for adaptive meshing on each of these problems. Additional experiments and sensitivity analysis is provide in the Appendix D. Full code to build the datasets and reproduce our results can be found at https://github.com/JRowbottomGit/g-adaptivity.
@@ 5.1 Experimental details
Our experimental pipeline consists of three parts: (i) we build datasets containing information about the PDE, FEM solution on a regular (i.e. not relocated) grid and the corresponding approximation of the Hessian of the solution; (ii) then we train either our model or the baseline to predict a relocated mesh on which we perform another FE solve to obtain the improved solution approximation; (iii) finally, we compare this FEM solution to a reference solution (calculated on a fine reference mesh) and determine the change in $L^{2}$-error over the original undeformed mesh, i.e. in the above notation we look at relative error reduction of $E(U_{\mathcal{M}_{\theta}})$ over $E(U_{\mathcal{Z}_{0}})$. These steps are repeated for the three PDE datasets as described below. Note that the Firedrake adjoint solve is only required during training of our G-Adaptive network and not during inference. The times reported in the below numerical examples thus contain a true reflection of the fast online mesh adaption times.
@@
For each experiment described below we build a randomised dataset with held-out test data. This means in each case we specify a set of solution values through varying source terms or boundary conditions (adjusted to the PDE at hand). We then generate training and test sets of coarse, undeformed, meshes $\mathcal{T}$ of varying resolution with associated FEM solution values and, for each mesh, we also compute a reference solution (on a finer reference mesh) which serves as comparison for the error computation. For most test cases the solution values are generated by randomly sampling Gaussians in the domain $\Omega$. For the example of the Navier–Stokes flow around a cylinder we used snapshots of a time-series simulation of vortex shedding. Full details on the specific configurations for each experiment are provided in Appendix C.
@@
Here we refer to our framework G-adaptivity and our model Diffformer synonymously, which we train using the regularized PDE-loss (9) (which is the regularized $L^{2}$ FEM approximation error + equidistribution regularizer). Calculation of the $L^{2}$-error is obtained by calculating an FE solution on the moved mesh and comparing this to the projection of the reference solution of a fine regular reference mesh onto sufficiently higher order elements. We train our new model (G-Adapt) and UM2N-G for 300 epochs using an Adam optimiser and learning rate of 0.001. Our model has $4$ diffformer blocks as described in Appendix C. Each blocked is rolled out using explicit Euler integration for 32 steps with a step size of 0.1. For the baseline UM2N we trained using 1000 epochs in order to achieve good performance but we believe further tuning of the training may be required in order to achieve a similar performance to the one reported in (Zhang et al., 2024). UM2N remains an important baseline and we expect that with appropriate training the method would be able to achieve a similar error reduction (ER) as the Monge-Ampère (MA) solver, but we would like to highlight that even in the best reported results of the original paper, UM2N never achieved a larger error reduction than MA.
@@
We report three metrics to evaluate the performance of the mesh relocation methods at hand: (i) the relative $L^{2}$ error reduction (ER) of the FE solution on the relocated mesh versus the FE solution on the initial coarse mesh (larger error reduction means improved performance); (ii) the time taken to relocate the mesh (shorter times means faster relocation); (iii) the aspect ratio of the deformed mesh as a measure of mesh quality as described in Appendix F.4 (a single digit aspect ratio is generally acceptable and a smaller aspect ratio indicates a more regular mesh). Each experiment is performed in full five times (training and evaluation) with different random seeds to provide the error bars.
@@ 5.2 Benchmarking on Poisson’s equation
Our first benchmark is on the classical Poisson problem $-\nabla^{2}u=f({\mathbf{z}})$ with Dirichlet boundary conditions. Full details of the FEM formulation are given in Appendix C.1. We benchmark the results against two datasets on a square (cf. Figure 1) and polygonal domain respectively (cf. Figure 4). Further evaluations on five additional (non-convex) geometries are provided in Appendix D.4. For both aforementioned examples we sample source terms and boundary conditions corresponding to underlying Gaussian fields. On the square domain $\Omega=[0,1]^{2}$ we initialise the mesh-deformation with a regular grid and to showcase G-adaptivty’s ability to work on irregular domains with unstructured meshes we apply a similar methodology to the convex polygonal dataset (a sample is shown in Figure 4). The results for both datasets are presented in Table 9. The central observation in these results is that our methodology provides the very first ML approach to mesh relocation which is able to outperform MA in terms of error reduction, while retaining the fast mesh relocation times given by the state-of-the-art GNN - UM2N (Zhang et al., 2024).
@@
$\dagger$ The direct optimization method is included here purely for exposition, showing that MA-meshes are not necessarily optimal. DirectOpt computes the optimal mesh for a given PDE with known solution but is extremely slow and relies on data which is not available during inference, thus it does not constitute a practical adaptive meshing strategy. In contrast, once trained, our G-Adaptive approach yields fast online mesh movement without needing reference solution values.
@@ 5.3 Time-dependent Burgers’ equation
We train the models on a set of Gaussian initial conditions for a timestep $\delta t=0.02$ with 2 steps and evaluate by following 10 trajectories of randomly sampled Gaussians in the Burgers equation for 20 timesteps, remeshing after every 2 steps (cf. Figure 5 and Appendix D.3). The results in the top part of Table 2 show the average error reduction over achieved over every block of two timesteps. While the MA performs well on this task, we note that the UM2N and UM2N-G baselines appear to lead to a negative error reduction (i.e. an increase), which is likely due to the fact that the Burgers’ equation changes the solution shape and thus trajectories will lead to out-of-distribution cases for methods that are trained only on initial conditions. Due to the structural regularity of our new approach (cf. Section 4.2) our approach is able to deal with out-of-distribution data very well, and most importantly is able to outperform MA in terms of error reduction while retaining a fast mesh relocation time.
@@ 5.4 Navier–Stokes equation and flow past a cylinder
5.5 3D adaptive meshing
The G-Adaptivity framework and diffusion deformer model are also easily adapted to the 3D setting. To demonstrate this we perform an experiment on a 10x10x10 unit cube for the 3D Poisson problem with Dirichlet boundary conditions and Gaussian solutions, analogous to Section 5.2. An example of the corresponding results can be seen in Figure 7. In the interest of brevity, the full numerical results are presented in Appendix 5.6 and show that the method outperforms MA significantly (out-of-the-box UM2N does not apply in 3D) and that it leads to effective mesh point concentration in regions of interest.
Figure 7: Examples of 3D solution fields and adapted meshes.
5.6 Scalability of the G-Adaptivity framework
The G-Adaptivity framework is able to scale to very large meshes. In particular the inductive learning property of GNNs ensures the ability of GNNs to transfer to unseen graphs in this case meaning we can perform super-resolution. In Table 9 we report experiments where the model is trained on 15x15 mesh and inference is performed on larger 60x60 (3,600 nodes) and 150x150 (22,500 nodes) meshes for the Poisson problem with 128 sampled Gaussians (see Figure 8). In order to scale the transformer encoder, which in naive form scales with $\mathcal{O}(N^{2})$ edges we use a sliding window SWIN (Liu et al., 2021) style transformer to capture the monitor function embedding at the mid-length scales. Our model consistently achieved significant mesh adaptation, accuracy improvement, and computational acceleration compared to Monge-Ampére, matching the performance observed on smaller-scale experiments.
Figure 8: G-Adaptivity on large-scale fine meshes
@@ 6 Conclusions and future work
Finally, we note that any machine learning-based approach is inherently statistical in nature, meaning that GNN-based meshing tools are likely to perform worse on out-of-distribution data. We observed this in our experiments with both pre-trained UM2N models and our own G-Adaptive approach when applied to PDEs whose solutions exhibited markedly different scales and features from those seen during training. Enhancing the scale-generalisation capabilities of ML-based adaptive meshing therefore remains an important open problem for future investigation.
@@ Acknowledgements
The authors would like to thank Patrick Farrell and David Ham for helpful advice related to capabilities of Firedrake, as well as Joseph G. Wallwork and Mingrui Zhang for sharing code and advice on the use of the UM2N baseline model and the Python Movement package. JR, GM, TD, CBS & CB gratefully acknowledge support from the EPSRC programme grant in ‘The Mathematics of Deep Learning’, under the project EP/V026259/1. GM gratefully acknowledges funding from the Mathematical Institute, University of Oxford. KS gratefully acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 850941).
@@ A.3 Gaussian basis function expansion for time-dynamic training
Appendix B Notes on the use of Firedrake in G-Adaptivity
@@
B.1 Additional details on implementation
Training the GNN requires computing the derivative of the loss function $E(Z,U_{Z})$ with respect to node coordinates $Z$. Since $E(Z,U_{Z})$ is a PDE-constrained functional, it is necessary to use adjoint models to compute these derivatives efficiently. The derivative and adjoint formulas depend on the loss function and its PDE constraints and automating their derivation is crucial to develop a general $r$-adaptivity methodology that can be trained seamlessly on different test cases. Firedrake is the perfect tool for this because it can derive adjoint models (Farrell et al., 2013; Mitusch et al., 2019) and automatically compute derivatives of $E(Z,U_{Z})$ with respect to node coordinates (Ham et al., 2019b). Deriving these formulas by hand is nontrivial, tedious, and error prone. Firedrake is fully integrated with PyTorch (Bouziani et al., 2024), and this is key to formulate hybrid FEM-torch architectures required to train the GNN. Implementing our approach in Firedrake required minimal adaptations: the GNN model must conform to the Firedrake external operator API, and a term must be added to the derivatives with respect to node coordinates when $E(Z,U_{Z})$ comprises a finite element solution computed on a finer mesh.
As an example consider the shape derivative $dJ(Z,U_{Z})[T]$ of the functional $J(Z,U_{Z})=||U_{Z}||^{2}_{L_{2}(\Omega)}$, which is a simplified version of $E(Z,U_{Z})$ in (3). The constraint on $U_{Z}$ is given by the simplest testcase: the weak Poisson equation in Appendix C.1 with $f=4$. With Firedrake and PyAdjoint, we can compute $dJ(Z,U_{Z})[T]$ as follows:
mesh = UnitSquareMesh(3, 3)
continue_annotation()
Q = mesh.coordinates.function_space()
T = Function(mesh.coordinates.function_space())
mesh.coordinates.assign(mesh.coordinates + T)
V = FunctionSpace(mesh, CG, 1)
u = Function(V)
v = TestFunction(V)
solve((dot(grad(u),grad(v))-4*v)*dx==0, u, bcs=DirichletBC(V, 0, on_boundary))
J = assemble(u**2*dx)
Jred = ReducedFunctional(J, Control(T))
Jred.derivative()
Crucially, this only requires the implementation of the forward constraint equation in Appendix C.1. On the other hand, a tedious manual derivation of $dJ(Z,U_{Z})[T]$ leads to $dJ(Z,U_{Z})[T]=\int_{\Omega}(U_{Z}^{2}+\nabla U_{Z}\cdot\nabla p-4p)\nabla\cdot T-\nabla U_{Z}(DT+DT^{\top})\nabla p\;dx$ with $p$ being the (weak) solution of the adjoint equation $\Delta p=2U_{Z}$. These formulae are problem dependent and will be significantly more complicated for other PDE constraints. For test cases such as the Navier Stokes equations in Appendix C.3 this approach quickly becomes intractable, as highlighted in (Ham et al., 2019b, p. 1818).
In contrast, adapting the code above to the problems described in Appendix C.2 & C.3 requires only minor changes.
Appendix C Mathematical description of the numerical experiments
@@ D.1 Model and data hyperparameters
D.2 Additional results in the Poisson square and polygon case
In Figures 9 and 10 we present additional plots from the Poisson experiments on the square and polygonal domain detailed in the main paper exhibiting the types of meshes generated with our novel G-Adaptive methodology.
@@ D.2 Additional results in the Poisson square and polygon case
D.3 Additional examples of Burger’s evolution
In Figure 11 we include some additional mesh trajectories from the experiment performed in Section 5.3.
Figure 11: Snapshots of the velocity field (x-component) together with the corresponding deformed meshes provided by Monge–Ampère.
D.4 Further experiments on non-convex domains
While section 5.4 already contains an example of a non-convex domain we provide further evidence that our method extends to this case using domain data from UM2N (Zhang et al., 2024). In particular we conducted experiments similar to the setup of 5.2 on five non-convex domains, cf. Figure 12. On each domain we solve Poisson’s equation for randomly sampled Gaussian solutions with 100 training datapoints and 100 unseen test datapoints.
The results (error reduction scores are listed in in Table 5 and the full results can be seen in Figure 12) confirm that our method performs robustly on non-convex geometries, achieving significantly greater error reduction than baselines and generating regular non-tangled meshes on all tested domains, succeeding even when some other approaches fail. Note that the UM2N results reported below were obtained using the pretrained model from the UM2N repository, since the MA meshes obtained using (Wallwork et al., 2024) were unsuitable for direct training in these cases.
Table 5: Error reduction scores (%) on Poisson’s equation in non-convex domain.
[…]
Geometry 1 0.23 $\pm$ 0.00 $-76.85\pm 0.00$ 1.92 $\pm$ 0.02 7.97 $\pm$ 0.04
Geometry 2 $-1.00\pm 0.00$ $-83.88\pm 0.00$ 0.69 $\pm$ 0.08 8.88 $\pm$ 0.24
Geometry 3 $-75.82\pm 0.00$ $-0.96\pm 0.04$ 6.62 $\pm$ 0.09
H-Geometry $-108.31\pm 0.00$ $-73.59\pm 0.00$ $-0.92\pm 0.00$ 7.51 $\pm$ 0.00
L-Geometry $-89.40\pm 0.00$ $-138.43\pm 0.00$ 13.94 $\pm$ 1.18 16.25 $\pm$ 0.25
* Monge–Ampère solvers in general struggle with non-convex domains. $\dagger$ Since the MA data available is not suitable for accurate training we use the pretrained model from (Zhang et al., 2024) for our evaluation.
Figure 12: Benchmarking results for Poisson’s equation on non-convex geometries.
D.5 Model hyper-parameter sensitivity analysis
We have performed extensive sensitivity studies and found that our approach is robust to the particular choice of hyperparameters for the diffformer blocks. We used $N_{b}=4$ blocks and rollout using explicit Euler time integration for 32 timesteps with a step size of 0.1. It should be noted that the hyperparameters were identical in all experiments performed in the paper and did not require finetuning to the specific problem. Tables 6 and 7 show the sensitivity of the model to the diffusion parameters in terms of error reduction and inference time.
Table 6: The effect of pseudotime and diffusion timesteps on the Error reduction.
Diffusion Time and time-step performance sensitivity
[…]
0.05 10.41 16.60 18.95 15.85 21.82 22.93
0.1 12.97 14.52 20.68 15.71 20.27 20.85
0.25 19.54 19.11 22.30 19.94 23.11 22.09
0.5 20.43 22.65 22.14 22.42 21.32 21.92
1 20.60 21.57 21.16 22.10 19.71 19.40
Table 7: The effect of pseudotime and diffusion timesteps on inference time (ms).
Diffusion Time and time-step inference time
[…]
0.05 60 44 46 116 65 247
0.1 54 42 79 61 86 208
0.25 41 40 48 56 119 108
0.5 50 58 49 92 125 158
1 52 59 45 69 100 108
We investigated the sensitivity of G-Adaptivity to the weighting of the equidistribution regularizer in Section 4.4. We found an optimal value in the Poisson square example was a loss weighting of 1, which was used throughout all experiments.
Table 8: The effect of equidistribution loss regularisation on Error reduction.
Equidistribution loss regularizer performance sensitivty
Reg. weight 0 (no equi.-dist. loss) 0.5 1 2 4 8
Error Red. (%) 22.42 22.95 23.99 23.21 22.14 20.96
D.6 Scalability of the G-Adaptivity framework
In Section 5.6 we claimed the G-Adaptivity framework is able to scale to very large meshes via super-resolution. In Table 9 we report experiments where the model is trained on 15x15 mesh and inference is performed on larger 60x60 (3,600 nodes) and 150x150 (22,500 nodes) meshes for the Poisson problem with 128 sampled Gaussians (see Figure 8). We also show the results for G-Adaptivity applied to a 10x10x10 (1,000 node) cube (see Figure 7).
Table 9: Benchmarking results of scaling G-Adaptivity on Poisson Square and Cube datasets.
Scaling G-Adaptivity
Scale Model Error Red. (%) Time (ms)
60x60 MA 11.94 ± 1.50 23084
G-Adapt 27.47 ± 0.89 452
150x150 MA 17.96 ± 115395
G-Adapt 25.70 ± 1.51 2555
10x10x10 MA 12.71 ± 0.00 41049
G-Adapt 28.08 ± 0.36 494
@@ Appendix E Further details on mesh relocation
E.1 Advantages and limitations of $r$-adaptivity
$r$-adaptivity is a newer technology than $h$-adaptivity and as such is not yet widely adapted in industry. However, it has certain significant advantages over $h$-adaptivity. In particular it works with a constant data structure, is easy to use on parallel architectures, it gives a more regular mesh (often with guaranteed mesh regularity), it naturally inherits Lagrangian and scaling structures in a PDE (which is very useful for example in ocean modelling and studying PDEs with singularities), and can be easily linked to existing external software designed to solve a PDE on an unstructured mesh (for example a discontinuous Galerkin solver). As a result, $r$-adaptive methods have recently been very successfully used, for example, in the operational data assimilation codes of national weather forecasting offices, which when coupled to the computational dynamical core, have led to a very significant increases in computational accuracy, particularly for resolving local weather features such as fog and ice (Piccolo & Cullen, 2012). $r$-adaptivity has also found natural applications in the steel industry where the Lagrangian nature of the approach is very well suited to the evolving fine structures in the forging process (Uribe et al., 2024). Possible disadvantages of $r$-adaptivity, such as excessive mesh computation cost, and a tendency to mesh tangling, are exactly the issues we address in this paper, proposing a fast and accurate method which avoids tangling.
E.2 The Equidistribution Principle
@@
where $\theta$ (to be determined) is a constant. The function $m$ is usually a function of $u$. An important example is given by the problem of linearly interpolating $u({\mathbf{z}})$ as it follows from Céa’s lemma that the resulting interpolation error is an (often tight) upper bound for the FE solution error. In this case $m$ will be a function of the curvature of $u$ (with the exact form dependent on the norm used to measure the error) (Huang & Russell, 2011).
@@
In one dimension the equidistribution condition (24) uniquely defines each cell length, and thus the cell shape, and hence the whole mesh. However in two dimensions it only gives the cell area but not the shape. To find the mesh uniquely additional conditions must be imposed. Noting the correspondence between the equidistribution condition and a measure on the physical space, the deformation map can be viewed as mapping a uniform measure in the computational space to a new measure in the physical space. It is natural to seek a map which minimises the cost of doing this, as this leads to meshes in the physical domain which are close to uniform and hence have minimal skewness and which avoid tangling. This gives an obvious link between mesh generation and optimal transport. In the continuous setting such a map can be calculated by solving (either directly or by using a surrogate solver) an associated Monge-Ampére equation, leading to the MA methods described in the main body of the text. Note that with modifications this procedure can also be used to generate meshes on non-planar manifolds (McRae et al., 2018).